OpenStudy (anonymous):

find the derivative of (x)/(x+5)

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=x, f'(x)=1, g(x)=x+5,g'(x)=1\]

OpenStudy (anonymous):

i couldn't read your reply

zepdrix (zepdrix):

Does the fancy text not show up? Are you using Internet Explorer by chance? :o

OpenStudy (anonymous):

yes and yes

zepdrix (zepdrix):

Yah Internet Explorer doesn't play nicely with the LaTeX Plugin :( You should try another browser. If you don't feel like doing that, I can try to explain it in regular text. c:

OpenStudy (anonymous):

yes could you please I used [g(x)*f'(x) - f(x)* g'(x)] / (g(x))^2 --the product rule is that right

zepdrix (zepdrix):

Yes, your quotient rule looks correct. So we get something like this: [(x+5)*(x)' - (x)*(x+5)'] / (x+5)^2 Right? Where the primes are the terms we need to differentiate still.

OpenStudy (anonymous):

i think i did something wrong i got [(x+5)*(1)-(x)*(1)] / (x+5)^2

OpenStudy (espex):

How about if you re-wrote it and used the product rule. \[x(x+5)^{-1}\]

zepdrix (zepdrix):

Yes very good Mizz, see how I had primes on the (x)' and (x+5)' terms? That let's us know that we need to take the derivative of those terms. I was simply doing the set up. The next step would give you what you actually came up with.